# U-Motion: Learned Point Cloud Video Compression with U-Structured Temporal Context Generation

### v2

@@ 3.1 Overview of Attribute Compression
As shown in Figure 1, U-Motion is a hierarchical architecture consisting of $K+1$ ($K=4$ during our experiments) U-Inter and Spatial Predictive Coding layers, where the $l$-th layer corresponds to the $l$-th level of the point cloud when represented by an octree, with the highest layer $H$ corresponding to the original voxelized point cloud. We denote the point cloud frame (including both geometry coordinate and color attribute for each point at each octree level) at time $t$ and layer $l$ as $p_{l}^{t}$, and its reference (i.e. the decoded previous point cloud frame) as $\hat{p}_{l}^{t-1}$. $p_{l}^{t}$ is expressed as a sparse tensor $\left[G(p_{l}^{t}),A(p_{l}^{t})\right]$, where $G(p_{l}^{t})$ denotes an $N\times 3$ geometry coordinate matrix and $A(p_{l}^{t})$ denotes the corresponding $N_{l}\times 3$ attribute matrix, with $N_{l}$ denoting the total number of points at layer $l$. The encoder fuses $p_{l}^{t}$ and the upper layer’s latent $f_{l+1}^{t}$ (after down-sample) to generate the latent representation $f_{l}^{t}$. The same encoder is also used to generate the reference frame’s latent $\hat{f}_{l}^{t-1}$.
@@ 3.3 U-Inter
Instead of compressing $\tilde{e}_{l}^{t}$ directly, we propose a motion propagation block that fuses $\tilde{e}_{l}^{t}$ with $\tilde{e}_{l+1}^{t}$ (through a top-down connection), and generates fused motion feature $e_{l}^{t}$, which contains motion information of layers $l:H$. This block enables 1) propagation of high-frequency details to the coarser scale, important for correct motion estimation at each scale; 2) distribution of finer-scale motion across multiple scales, to facilitate more efficient and flexible compression of motion information. In later ablation studies, we will show that when this propagation block is removed, the network often yields degraded motions in part due to the rate constraint. Note that this top-down motion propagation introduces redundancies among $e_{l}^{t}$ across layers. Therefore, the bottom-up motion prediction module contextually detaches the redundancy in $e_{l}^{t}$ that is predictable from the decoded coarser-scale $\hat{e}_{l-1}^{t}$, when compressing $e_{l}^{t}$, and contextually restores $\hat{e}_{l-1}^{t}$ into $\hat{e}_{l}^{t}$ on the decoder side.
@@
Although [17] introduces coarse-to-fine motion estimation, it only applies motion compensation at the finest scale to code the latent features using temporal prediction at the finest scale. We propose progressive coarse-to-fine motion compensation with multiples scales of compensation and predictive coding. The motion decoding module decodes a residual motion $\Delta m_{l}^{t}$ from motion feature $\hat{e}_{l}^{t}$, which is then added to the upsampled $\hat{m}_{l-1}^{t}$ to generate the motion $\hat{m}_{l}^{t}$.
@@
As shown in [21] for motion estimation in 2D video, different feature channels may experience different motions. Borrowing this idea, we estimate multiple motion groups at each scale. The motion decoding module outputs a motion $\hat{m}_{l}^{t}\in\mathbb{R}^{N_{l}^{t}\times G_{l}\times 3}$, where $N_{l}^{t}$ is the number of points in frame $t$ layer $l$, $G_{l}$ is the number of motion groups of layer $l$. Assuming the latent dimension is $C_{l}$, we divide the $C_{l}$ feature channels evenly into $G_{l}$ groups, and decode from the motion features $\hat{e}_{l}^{t}$ a residual motion flow for each group, $\Delta m_{l,g}^{t}$. In general, we find that increasing the number of groups brings gain, despite the potential motion bit rate increase brought by coding more motion groups. However, increasing the group size results in a quadratic increase in complexity for motion compensation. Consequently, using large motion groups at fine scales leads to a sharp rise in computational complexity. Therefore, we employ motion group merging, utilizing large group sizes at coarse scales and progressively merging and reducing the number of motion groups at finer scales.
@@ 2.1 Settings for Attribute Compression
Our U-Motion attribute compression model consists of five layers, i.e. five U-Inter layer and five Spatial Predictive Coding layer. We adopt the base layer strategy employed in YOGA [35], which compresses the lower-scale point cloud $p_{H-3}^{t}$ with G-PCC-GesTM to provide a base for the hierarchical reconstruction. Note that for layer $l\in\left[L,H-3\right]$, although the color is coded by GesTM, there still exists the motion embedding $e_{l}^{t}$ and latent residual $r_{l}^{t}$ to be coded by our network for higher layers’ reconstruction. We do not losslessly compress $p_{l}^{t}$ at layer $H-3$ due to the excessive bit-rate consumption. Instead, we use different GesTM quantization parameters ($QP$) for different rate points to achieve the R-D balance of overall performance. We traverse different combinations of $\lambda$ and $QP$ on a subset of dataset and get the most R-D efficient pairs: $\lambda=$[300, 460, 705, 910, 1655, 2537, 3888, 5960, 9134, 14000, 16000], and the corresponding $QP=$[20, 20, 16, 16, 16, 16, 12, 12, 8, 8, 8].

### v4

@@ 3.1 Overview of Attribute Compression
As shown in Figure 1, U-Motion is a hierarchical architecture consisting of $K+1$ ($K=4$ during our experiments) U-Inter and Spatial-Temporal Predictive Coding layers, where the $l$-th layer corresponds to the $l$-th level of the point cloud when represented by an octree, with the highest layer $H$ corresponding to the original voxelized point cloud. We denote the point cloud frame (including both geometry coordinate and color attribute for each point at each octree level) at time $t$ and layer $l$ as $p_{l}^{t}$, and its reference (i.e. the decoded previous point cloud frame) as $\hat{p}_{l}^{t-1}$. $p_{l}^{t}$ is expressed as a sparse tensor $\left[G(p_{l}^{t}),A(p_{l}^{t})\right]$, where $G(p_{l}^{t})$ denotes an $N\times 3$ geometry coordinate matrix and $A(p_{l}^{t})$ denotes the corresponding $N_{l}\times 3$ attribute matrix, with $N_{l}$ denoting the total number of points at layer $l$. The encoder fuses $p_{l}^{t}$ and the upper layer’s latent $f_{l+1}^{t}$ (after down-sample) to generate the latent representation $f_{l}^{t}$. The same encoder is also used to generate the reference frame’s latent $\hat{f}_{l}^{t-1}$.
@@ 3.3 U-Inter
Instead of compressing $\tilde{e}_{l}^{t}$ directly, we propose a motion propagation block that fuses $\tilde{e}_{l}^{t}$ with $e_{l+1}^{t}$ (through a top-down connection), and generates fused motion feature $e_{l}^{t}$, which contains motion information of layers $l:H$. This block enables 1) propagation of high-frequency details to the coarser scale, important for correct motion estimation at each scale; 2) distribution of finer-scale motion across multiple scales, to facilitate more efficient and flexible compression of motion information. In later ablation studies, we will show that when this propagation block is removed, the network often yields degraded motions in part due to the rate constraint. Note that this top-down motion propagation introduces redundancies among $e_{l}^{t}$ across layers. Therefore, the bottom-up motion prediction module contextually detaches the redundancy in $e_{l}^{t}$ that is predictable from the decoded coarser-scale $\hat{e}_{l-1}^{t}$, when compressing $e_{l}^{t}$, and contextually restores $\hat{e}_{l-1}^{t}$ into $\hat{e}_{l}^{t}$ on the decoder side.
@@
Although [17] introduces coarse-to-fine motion estimation, it only applies motion compensation at the finest scale to code the latent features using temporal prediction at the finest scale. We propose progressive coarse-to-fine motion compensation with multiples scales of compensation and predictive coding. The motion decoding module decodes a residual motion $\Delta m_{l}^{t}$ from motion feature $\hat{e}_{l}^{t}$, which is then added to the upsampled $m_{l-1}^{t}$ to generate the motion $m_{l}^{t}$.
@@
As shown in [21] for motion estimation in 2D video, different feature channels may experience different motions. Borrowing this idea, we estimate multiple motion groups at each scale. The motion decoding module outputs a motion $m_{l}^{t}\in\mathbb{R}^{N_{l}^{t}\times G_{l}\times 3}$, where $N_{l}^{t}$ is the number of points in frame $t$ layer $l$, $G_{l}$ is the number of motion groups of layer $l$. Assuming the latent dimension is $C_{l}$, we divide the $C_{l}$ feature channels evenly into $G_{l}$ groups, and decode from the motion features $\hat{e}_{l}^{t}$ a residual motion flow for each group, $\Delta m_{l,g}^{t}$. In general, we find that increasing the number of groups brings gain, despite the potential motion bit rate increase brought by coding more motion groups. However, increasing the group size results in a quadratic increase in complexity for motion compensation. Consequently, using large motion groups at fine scales leads to a sharp rise in computational complexity. Therefore, we employ motion group merging, utilizing large group sizes at coarse scales and progressively merging and reducing the number of motion groups at finer scales.
@@ 2.1 Settings for Attribute Compression
Our U-Motion attribute compression model consists of five layers, i.e. five U-Inter layer and five Spatial-Temporal Predictive Coding layer. We adopt the base layer strategy employed in YOGA [35], which compresses the lower-scale point cloud $p_{H-3}^{t}$ with G-PCC-GesTM to provide a base for the hierarchical reconstruction. Note that for layer $l\in\left[L,H-3\right]$, although the color is coded by GesTM, there still exists the motion embedding $e_{l}^{t}$ and latent residual $r_{l}^{t}$ to be coded by our network for higher layers’ reconstruction. We do not losslessly compress $p_{l}^{t}$ at layer $H-3$ due to the excessive bit-rate consumption. Instead, we use different GesTM quantization parameters ($QP$) for different rate points to achieve the R-D balance of overall performance. We traverse different combinations of $\lambda$ and $QP$ on a subset of dataset and get the most R-D efficient pairs: $\lambda=$[300, 460, 705, 910, 1655, 2537, 3888, 5960, 9134, 14000, 16000], and the corresponding $QP=$[20, 20, 16, 16, 16, 16, 12, 12, 8, 8, 8].